Many expressions in special relativity contain the gamma factor Number where v is the speed of some object or reference frame, and c is the speed of light. Find the numerical value of for each of the speeds listed at right. Below, select the graph that most accurately represents how the 055cY- Number value of y depends on speed. Number
Answer #1

Gamma factor  \gamma=\frac{1}{\sqrt{1-v^2/c^2}}

For v=0 , \gamma=\frac{1}{\sqrt{1-(0)^2/c^2}}=1.00

For v=0.55c , \gamma=\frac{1}{\sqrt{1-(0.55 c)^2/c^2}}=1.20

For v=0.98c , \gamma=\frac{1}{\sqrt{1-(0.98 c)^2/c^2}}=5.02

The correct representation of graph for the dependence of \gamma on speed v is